# Polynomial Potential Inflation in the ACT Era: From CMB to Primordial Black Holes

### v2

@@ II.1 $\mathbf{n=2}$
For hill-top inflation, the values of scalar field at the end of inflation $\phi_{e}$ is determined by the condition
@@
we then get
@@ II.2 $\mathbf{n=3}$
Note that $\xi=0$ for the case of $n=2$. In terms of Eq. (II) and Eqs. (12)-(14), The running of the spectral index $\alpha_{s}$ can be written as
@@ II.3 $\mathbf{n=4}$
These relations allow us to map observational constraints directly onto the parameter space of the model. For the quartic potential, the cubic equation $V_{\phi}=0$ determines the critical points. Assuming one real root $\gamma$ and a pair of complex conjugate roots $\alpha\pm i\beta$, the integral can be performed analytically via partial fraction decomposition, giving
@@ III.1 Primordial curvature perturbations
We can set $a_{2}=0$ through a transformation $\phi\to\phi+\text{const}$ without loss of generality. Since we are considering small-field inflation, where the inflaton rolls down from near $\phi=0$, we can neglect terms higher than the third order when $\phi$ is in the vicinity of the origin
@@
The slow-roll parameters are
@@
$N_{e}=\int^{\phi_{e}}_{\phi_{*}}\frac{\mathrm{d}\phi}{\sqrt{2\epsilon}}=\left.\frac{1}{\sqrt{3a_{1}a_{3}}}\arctan\left(\sqrt{\frac{3a_{3}}{a_{1}}}\phi\right)\right|_{\phi_{*}}^{\phi_{e}}\,.$ (26)
Introducing a new quantity $N_{\mathrm{tot}}$ which characterize the total number of e-folding during inflation Baumann et al. (2008); Linde and Westphal (2008)
@@
$n_{s}-1\simeq 2\eta(\phi_{*})$ (28)
Combining Eqs. (26), (27) and (28), we obtain
$n_{s}-1=\frac{2\pi}{N_{\mathrm{tot}}}\tan\left(\frac{\pi N_{e}}{2N_{\mathrm{tot}}}-\arctan\left(\frac{N_{\mathrm{tot}}}{2\pi}\right)\right)\,.$ (29)
@@
$n_{s}-1=-\frac{2\pi}{N_{\mathrm{tot}}}\cot\left(\frac{\pi N_{e}}{2N_{\mathrm{tot}}}\right)\,.$ (30)
To coincide with the CMB observation, we need a red-tilted power spectrum, which means $N_{\mathrm{tot}}\geq N_{e}$. If we set $N_{e}=30$ and assume $a_{3}\sim\mathcal{O}(1)$,then we have
@@
$\alpha_{s}=\frac{dn_{s}}{d\ln k}=-\frac{\pi^{2}}{N_{\mathrm{tot}}^{2}}\csc^{2}\left(\frac{\pi N_{e}}{2N_{\mathrm{tot}}}\right)\,.$ (33)
@@
In the vicinity of this point, the potential becomes extremely flat, causing the inflaton to enter a phase of ultra-slow-roll inflation. During this phase, the second slow-roll parameter $\eta$ becomes large and negative, leading to an exponential growth of the curvature perturbations. The location of the inflection point determines the comoving wavenumber $k_{p}$ at which the power spectrum peaks, and the flatness of the plateau controls the amplification factor.
As a concrete example, we obtain a viable power spectrum consistent with CMB constraints on large scales and a pronounced peak on small scales with the following parameter set
@@ III.2 Primordial black holes
PBHs are hypothesized to form in the early universe through the gravitational collapse of large‑amplitude curvature perturbations generated during inflation. When these perturbations re-enter the Hubble horizon in the radiation-dominated era, those exceeding a critical density threshold $\delta_{c}\approx 0.45$ can overcome radiation pressure and collapse into PBHs Musco et al. (2005, 2009); Musco and Miller (2013). This process requires a significant enhancement of the primordial curvature power spectrum $\mathcal{P}_{\mathcal{R}}(k)$ to $\mathcal{O}(10^{-2})$ on small scales ($k\gg 1~\mathrm{Mpc}^{-1}$), contrasting with the CMB scale value of $\mathcal{O}(10^{-9})$.
@@ IV Summary and discussion
For the quadratic potential ($n=2$), we find that the model can accommodate the ACT data at $1\sigma$ level within a limited range of the dimensionless coefficients $a_{1}$ and $a_{2}$. However, when imposing the additional constraint on the number of $e$-folds NN, the model struggles to simultaneously satisfy the observational bounds and the requirement of sufficient inflation ($N\gtrsim 60$).
@@
Our results highlight the importance of revisiting and refining inflationary models in light of the latest observational data. The polynomial potential, despite its simplicity, can still provide a viable framework for explaining the early Universe, particularly when extended to higher-order terms. However, the model’s parameter space is increasingly constrained by the precise measurements of the CMB anisotropies. To further test the polynomial potential inflation and other inflationary models, future CMB experiments with improved sensitivity and resolution, such as CMB-S4 Abazajian et al. (2016), Simons Observatory Ade et al. (2019), and LiteBIRD Allys et al. (2023), will be crucial. These experiments are expected to provide even tighter constraints on $n_{s}$, $r$, and $\alpha_{s}$, allowing for a more stringent model selection and parameter estimation.
@@ Acknowledgements.
This work has been supported by the National Key Research and Development Program of China (No. 2023YFC2206704). ZZP is supported by the National Natural Science Foundation of China under Grant No. 12447166. ZCC is supported by the National Natural Science Foundation of China under Grant No. 12405056, the Natural Science Foundation of Hunan Province under Grant No. 2025JJ40006, and the Innovative Research Group of Hunan Province under Grant No. 2024JJ1006. LL is supported by the National Natural Science Foundation of China (Grant No. 12505054 and  12433001).

### v3

@@ II.1 $\mathbf{n=2}$
The magnitude $|a_{1}|=\sqrt{r}/(2\sqrt{2})$ follows directly from $r=16\epsilon=8a_{1}^{2}$. The negative sign is then fixed by requiring $V_{\phi}(\phi_{*})<0$, ensuring that the inflaton rolls downhill in the positive field direction consistent with our hill-top configuration. For hill-top inflation, the values of scalar field at the end of inflation $\phi_{e}$ is determined by the condition
@@
This condition marks the breakdown of the slow-roll approximation and the end of the inflationary phase. It should not be confused with the true vacuum state ($V_{\phi}=0$), which the inflaton reaches only later during the reheating epoch. For hill-top models, $|\eta|=1$ is typically reached before $\epsilon=1$ because the potential curvature dominates over the slope near the maximum Liddle and Lyth (2000); Baumann (2011); Boubekeur and Lyth (2005). We then get
@@ II.2 $\mathbf{n=3}$
Here $k$ denotes the comoving wavenumber of primordial perturbations. The scalar spectral index $n_{s}$ characterizes the scale dependence of the primordial power spectrum, $\mathcal{P}_{s}(k)\propto k^{n_{s}-1}$, arising because different $k$-modes exit the Hubble horizon at different epochs during inflation and thus probe slightly different values of the slow-roll parameters. Observational constraints are quoted at the pivot scale $k_{*}=0.05~\mathrm{Mpc}^{-1}$. Note that $\xi=0$ for the case of $n=2$. In terms of Eq. (II) and Eqs. (12)-(14), The running of the spectral index $\alpha_{s}$ can be written as
@@ II.3 $\mathbf{n=4}$
These relations allow us to map observational constraints directly onto the parameter space of the model. For the quartic potential, the cubic equation $V_{\phi}=0$ determines the critical points. The critical points of the potential, determined by $V_{\phi}=0$, characterize the structure of the scalar potential (its maxima and minima). This is distinct from the end-of-inflation condition $|\eta(\phi_{e})|=1$, which identifies where slow-roll violation terminates the accelerated expansion. Assuming one real root $\gamma$ and a pair of complex conjugate roots $\alpha\pm i\beta$, the integral can be performed analytically via partial fraction decomposition, giving
@@ III.1 Primordial curvature perturbations
For the quintic potential ($n=5$), we adopt a different computational scheme from the cases $n=2$ to $n=4$ treated in Section II. Here, the field translation $\phi\to\phi+C$ is used to set $a_{2}=0$, simplifying the subsequent analysis. Unlike the previous cases where we set $\phi_{*}=0$ exactly, here the initial field value $\phi_{*}$ is not fixed to zero but is assumed to be small. Since the inflaton rolls down from near $\phi=0$, we can neglect terms higher than the third order when $\phi$ is in the vicinity of the origin
@@
For the analysis of large-scale observables (CMB/ACT), the field remains small, and we can safely neglect the higher-order terms $a_{4}\phi^{4}$ and $a_{5}\phi^{5}$. However, for SIGW and PBH production, the field approaches $\phi\sim\mathcal{O}(1)$ near the inflection point, and the quartic and quintic terms become important and must be retained. The slow-roll parameters are
@@
${N_{1}}=\int^{{\phi_{1}}}_{\phi_{*}}\frac{\mathrm{d}\phi}{\sqrt{2\epsilon}}=\left.\frac{1}{\sqrt{3a_{1}a_{3}}}\arctan\left(\sqrt{\frac{3a_{3}}{a_{1}}}\phi\right)\right|_{\phi_{*}}^{{\phi_{1}}}\,.$ (26)
Here $\phi_{1}$ and $N_{1}$ denote the field value and e-folding number at the end of the first slow-roll phase, not the end of inflation. In this phase, which corresponds to large-scale (CMB/ACT) observations, the field remains small: $\phi_{*}<\phi<\phi_{1}<1$ (in Planck units). This justifies neglecting higher-order terms in the polynomial potential. At smaller scales, when $\phi>\phi_{1}$ and the field approaches the inflection point, higher-order terms become significant and must be retained—particularly for analyzing enhanced perturbations relevant to PBH formation. Introducing a new quantity $N_{\mathrm{tot}}$ which characterize the total number of e-folding during inflation Baumann et al. (2008); Linde and Westphal (2008)
@@
$n_{s}-1\simeq 2\eta(\phi_{*}).$ (28)
For $N_{\mathrm{tot}}$ to be real and positive, the product $a_{1}a_{3}$ must be positive, requiring $a_{1}$ and $a_{3}$ to share the same sign. Meanwhile, the hill-top configuration demands $V_{\phi}=V_{0}(a_{1}+3a_{3}\phi^{2})<0$ during inflation, which is impossible if both $a_{1}$ and $a_{3}$ are positive. Therefore, both parameters must be negative: $a_{1}<0$ and $a_{3}<0$. This is consistent with the observed red-tilted spectrum, since $n_{s}-1\simeq 12\,a_{3}\,\phi_{*}<0$ for $a_{3}<0$ and $\phi_{*}>0$. Combining Eqs. (26), (27) and (28), we obtain
${n_{s}-1=\frac{2\pi}{N_{\mathrm{tot}}}\tan\left(\frac{\pi N_{1}}{2N_{\mathrm{tot}}}-\arctan\left(\frac{N_{\mathrm{tot}}}{\pi}\right)\right)\,.}$ (29)
@@
$n_{s}-1=-\frac{2\pi}{N_{\mathrm{tot}}}\cot\left(\frac{\pi{N_{1}}}{2N_{\mathrm{tot}}}\right)\,.$ (30)
To coincide with the CMB observation, we need a red-tilted power spectrum, which means $N_{\mathrm{tot}}\geq{N_{1}}$. If we set ${N_{1}}=30$ and assume $a_{3}\sim\mathcal{O}(1)$,then we have
@@
$\alpha_{s}=\frac{dn_{s}}{d\ln k}=-\frac{\pi^{2}}{N_{\mathrm{tot}}^{2}}\csc^{2}\left(\frac{\pi{N_{1}}}{2N_{\mathrm{tot}}}\right)\,.$ (33)
@@
In the vicinity of this point, the potential becomes extremely flat, causing the inflaton to enter a phase of ultra-slow-roll inflation. As the field evolves past the inflection point, the second slow-roll parameter $\eta$ transitions from positive to negative values. At the exact inflection point where $V_{\phi\phi}=0$, we have $\eta=0$. After crossing the inflection point, $\eta$ becomes increasingly negative, leading to the enhancement of primordial perturbations. The location of the inflection point determines the comoving wavenumber $k_{p}$ at which the power spectrum peaks, and the flatness of the plateau controls the amplification factor.
In the ultra-slow-roll regime near the inflection point, the field value $\phi$ is no longer small but approaches $\mathcal{O}(1)$. Consequently, the $a_{4}\phi^{4}$ and $a_{5}\phi^{5}$ terms, which were negligible for the large-scale analysis, now play a dominant role in determining the potential shape and the amplitude of enhanced perturbations. As a concrete example, we obtain a viable power spectrum consistent with CMB constraints on large scales and a pronounced peak on small scales with the following parameter set
@@ III.2 Primordial black holes
PBHs are hypothesized to form in the early universe through the gravitational collapse of large‑amplitude curvature perturbations generated during inflation. When these perturbations re-enter the Hubble horizon in the radiation-dominated era, those exceeding a critical density threshold $\delta_{c}$ can overcome radiation pressure and collapse into PBHs Musco et al. (2005, 2009); Musco and Miller (2013). This process requires a significant enhancement of the primordial curvature power spectrum $\mathcal{P}_{\mathcal{R}}(k)$ to $\mathcal{O}(10^{-2})$ on small scales ($k\gg 1~\mathrm{Mpc}^{-1}$), contrasting with the CMB scale value of $\mathcal{O}(10^{-9})$.
@@ IV Summary and discussion
For the quadratic potential ($n=2$), we find that the model can accommodate the ACT data at $1\sigma$ level within a limited range of the dimensionless coefficients $a_{1}$ and $a_{2}$. However, when imposing the additional constraint on the number of $e$-folds NN, the model struggles to accommodate the ACT data within the standard e-folding number range of $N\approx 50$$60$, as shown in Fig. 1. While larger values (e.g., $N=70$) can improve consistency, they are less typical based on thermal history considerations, highlighting the model’s limitations.
@@
For $n=2$ to $n=4$, our analytical results are exact within the slow-roll approximation with $\phi_{*}=0$. For $n=5$, our treatment relies on the small-field approximation valid during the first slow-roll phase ($\phi<\phi_{1}\ll 1$), which is appropriate for interpreting large-scale observations such as the ACT data. The analysis of small-scale dynamics near the inflection point, relevant for PBH production, would require including higher-order terms in the potential. Our results highlight the importance of revisiting and refining inflationary models in light of the latest observational data. The polynomial potential, despite its simplicity, can still provide a viable framework for explaining the early Universe, particularly when extended to higher-order terms. However, the model’s parameter space is increasingly constrained by the precise measurements of the CMB anisotropies. To further test the polynomial potential inflation and other inflationary models, future CMB experiments with improved sensitivity and resolution, such as CMB-S4 Abazajian et al. (2016), Simons Observatory Ade et al. (2019), and LiteBIRD Allys et al. (2023), will be crucial. These experiments are expected to provide even tighter constraints on $n_{s}$, $r$, and $\alpha_{s}$, allowing for a more stringent model selection and parameter estimation.
@@ Acknowledgements.
We thank the anonymous referee for the valuable and constructive suggestions. This work has been supported by the National Key Research and Development Program of China (No. 2023YFC2206704). ZZP is supported by the National Natural Science Foundation of China under Grant No. 12447166. ZCC is supported by the National Natural Science Foundation of China under Grant No. 12405056, the Natural Science Foundation of Hunan Province under Grant No. 2025JJ40006, and the Innovative Research Group of Hunan Province under Grant No. 2024JJ1006. LL is supported by the National Natural Science Foundation of China (Grant No. 12505054, 12447101 and  12433001) and the Fundamental Research Funds for the Central Universities.